The sum of two numbers is 41 and the difference is 5. What are the numbers?
i would like the answer but i would like to know how to do this

Answer by mananth(16946)

let one number be x
other be y .... x>y
x+y = 41
x-y =5
Add the two equations
2x=46
/2
x=23
ao y= 41-23=18
(23, 18)